# OmniGuard: An Efficient Approach for AI Safety Moderation Across Modalities

### v1

@@ 3.2.3 Setup for audio attacks
Following Section 3.1, we build a classifier that takes as input embeddings from layer 21 of Llama-Omni-8B. We train the classifier on the English portion of the training data we use for the text setting, by using a text-to-speech model to convert the text into audio. We use the open-source Kokoro model as the text-to-speech model.
@@
We are unaware of any existing models for detecting harmful audio input. The most relevant approach, SpeechGuard (Peri et al., 2024) adds noise as a defense against potentially harmful audio inputs but does not directly classify harmfulness. To contextualize our results for audio benchmarks, we compare performance to guard models that directly classify the raw text present in the audio (OmniGuard and LlamaGuard3).
@@
We use the two audio benchmarks (see Table 1 bottom row). We also evaluate on several text jailbreak benchmarks using Kokoro to convert them from text to speech: HB, FQ, Simple Safety Tests, SaladB, and TJS. We use Kokoro for generating text-to-speech versions.
@@ 4 Results
Table 3: Accuracy of detecting harmful prompts for text attack benchmarks that are (A) multilingual benchmarks, (B) English translated to 73 languages, and (C) English translated to languages not seen at training time. In all settings, OmniGuard achieves the highest performance. Table B1 further stratifies these results by high-resource, low-resource, and cipher languages.
@@
Table 3 compares the accuracy of detecting harmful prompts for text benchmarks. Table 3(A) shows results for multilingual benchmarks, where OmniGuard achieves the highest accuracy (86.36%) compared to the baselines, and achieves new state-of-the-art performance for 3 benchmarks: MultiJail, RTP-LX, and AyaRedTeaming. The strongest baseline is Polyguard, which yields an average accuracy of 83.19%, despite being trained on a much larger dataset (1.91M examples for Polyguard versus 103K examples for OmniGuard). In benchmarks that were translated from English to various other languages, including cipher languages, we again see that OmniGuard achieves the highest accuracy (Table 3(B)). Finally, Table 3(C) shows that OmniGuard outperforms finetuned versions of DuoGuard and Polyguard on unseen languages, demonstrating that OmniGuard can outperform methods that were trained specifically for multilingual harmfulness classification.
@@
Table 4 shows the accuracy of detecting harmful image and text prompts for (A) pairs consisting of images and text queries, where either the image or both the image and query can be harmful and (B) typographic images with various encryptions. OmniGuard achieves the highest performance for both sets of benchmarks (95.44% and 79.76%) while being trained using only about 3500 image-query pairs (compared to about 5500 datapoints used by LlavaGuard). The only benchmark where OmniGuard fails to detect harmful prompts is MML Base64, which consists of typographed images of prompts encrypted using base64 encoding.
@@
Table 5 shows the accuracy of detecting harmful audio prompts. OmniGuard detects harmful audio input with high accuracy across all benchmarks. As we are not aware of any existing defenses for audio jailbreaks, we compare against OmniGuard and LlamaGuard3’s accuracy in detecting harmful prompts when the same inputs are provided in English text. The accuracy OmniGuard achieves in detecting harmful audio inputs is similar to or higher than its performance for detecting harmful text inputs.
@@
We also evaluate the accuracy of OmniGuard and baselines in adapting to out-of-distribution code attacks given very few samples. In this setting, some prior work has speculated that guard models may be very data efficient, as they can make use of few-shot examples in-context (Inan et al., 2023). However, we find that baseline guard models generally struggle to rapidly adapt to this setting given few-shot examples (Figure 3). [^2] Note, we omit baseline guard models that achieve 90% accuracy or greater without any few-shot examples, as their training data likely explicitly includes code attacks. In contrast, OmniGuard is able to rapidly achieve close to 100% accuracy for all three benchmarks by updating its lightweight parameters using less than five examples.
@@ 5 Analysis
We perform ablation experiments to determine the effect of selecting the appropriate layer for training the OmniGuard classifier. For the text-only model, we compare the U-Score-selected layer (57) to 3 other layers (layer 10, layer 75, and the last layer) when used for a set of toxicity prediction tasks. Table 6 shows that the representations from the layer with the highest U-Score result in significantly better harmfulness classification accuracy, improving between 5% and 14% compared to the other layers.
@@
Table 6: OmniGuard’s accuracy of detecting harmful prompt when trained using representation from different model layers.
@@
OmniGuard is highly efficient at inference time because it re-uses the internal representations of the main LLM that is already processing the user query for generation. Therefore, its compute time is only that of a lightweight multilayer perceptron, making it much faster than baseline guard models (note that this does limit OmniGuard to only work when the generation model is open-source, so embeddings can be extracted). Table 7 shows the inference time required by various guard models to predict the harmfulness of prompts in the AdvBench dataset in English, translated to Spanish, French, Telugu, and base64 encoding. OmniGuard is the fastest and is about 120$\times$ faster than the fastest baseline (DuoGuard). Inference time as measured on a machine with 1 L40 GPU, 4 CPUs, and 50 GB RAM.
@@
Table 7: Average inference time required for harmfulness prediction on the AdvBench dataset (averaged over 5 languages). OmniGuard is about 120$\times$ faster than the fastest baseline (DuoGuard).
@@
We now analyze the harmfulness classification accuracy of OmniGuard by language, and compare it to the underlying LLM’s sentiment classification accuracy for the same language (Fig. 4). We measure harmfulness classification accuracy using OmniGuard on all the datasets in Table 3 and sentiment classification accuracy using Llama3.3-70B-Instruct with zero-shot prompting on 72 translated versions of the SST-2 dataset (translated to all the languages we consider).
We observe that the accuracies are generally correlated indicating that OmniGuard is able to defend well for languages for which the LLM is more coherent / susceptible to attack. Unsurprisingly, the accuracies for natural languages are higher than the accuracies for cipher languages. Nevertheless, harmfulness classification accuracy can be fairly high, even when sentiment classification accuracy is near chance (50%).
@@ 6 Related Work
Several techniques have recently emerged to attack or jailbreak LLMs. Early techniques relied on manual effort and were very time-intensive Shen et al. (2024b); Andriushchenko et al. (2024). Later techniques automated this process, e.g., Zou et al. (2023b); Jones et al. (2023); Zhu et al. (2023) proposed gradient-based approaches to identify inputs to jailbreak LLMs with white-box access. Another set of techniques start from a set of human written prompts and modify them using approaches like genetic algorithms (Liu et al., 2024a; Lapid et al., 2024; Li et al., 2024d), fuzzing (Yu et al., 2024a), or reinforcement learning (Chen et al., 2024b) to automatically produce prompts for jailbreaking. Another set of techniques, use a helper LLM to generate prompts that attack a target LLM (Chao et al., 2024; Ding et al., 2024; Mehrotra et al., 2024). Finally, Wei et al. (2024); Wang et al. (2023); Anil et al. (2024); Pernisi et al. (2024) use simple in-context demonstrations to jailbreak the models by overcoming its safety training and Russinovich et al. (2024); Li et al. (2024c) propose using multi-turn dialogues to jailbreak models.
@@
Most of the aforementioned jailbreak techniques focus on attacks in English, against which significant defense exists both at the model and system level. To tackle this, a novel set of techniques have emerged that attack models using inputs in various languages or obfuscations that are able to bypass the safety guardrails. Deng et al. (2024); Yong et al. (2024); Wang et al. (2024a); Yang et al. (2024); Yoo et al. (2024); Upadhayay and Behzadan (2024); Song et al. (2024) demonstrated that attacking models using mid and low resources languages led to higher attack success rates, compared to the case of attacking the model in high-resource languages like English.
Going beyond natural languages, a newer set of works propose using cipher characters or languages to evade the safety filters, e.g., Jin et al. (2024) propose interspersing cipher characters in between text, Jiang et al. (2024) propose replacing the unsafe words with their ASCII art versions, and Yuan et al. (2024a) propose prompting models in cipher languages like Morse, Atbash, Caesar.
@@
Using modalities apart from text aims to explore a completely new attack surface, like images or audios. Several recent works have shown that MLLMs remain vulnerable to being jailbroken when prompted with images or audios that have a harmful query (the same harmful query in text would be easily detected as harmful). Liu et al. (2024b) show that using a prompt with a correlated image, e.g., using an image of a bomb when asking the model to answer the question: How to make a bomb? is more likely to jailbreak a model than when using an uncorrelated image. Hu et al. (2025) argue that providing a harmful image with a benign query (see Figure 1) further increases the potential of jailbreaking the model. Gong et al. (2025) and Wang et al. (2024b) demonstrate jailbreaking models by simply typographically embedding harmful queries in an image.
@@
Safety moderation in LLMs broadly fits into two categories: intrinsic and extrinsic. Intrinsic mechanisms include finetuning or RLHF training on an LLM (Bianchi et al., 2024; Chen et al., 2024a; Yuan et al., 2024b; Ouyang et al., 2022b; Bai et al., 2022; Dai et al., 2023). Extrinsic safety mechanisms utilize external models to detect harmful inputs and responses; these models can either be simple filters or use guard models. Jain et al. (2023); Alon and Kamfonas (2023); Hu et al. (2024) propose using perplexity filtering for detecting harmful prompts. Guard models defend LLMs by training separate LLMs to detect harmful text (see Table 2).
@@
DuoGuard (Deng et al., 2025) and PolyGuard (Kumar et al., 2025) are the two previous guard models that were specifically trained to defend against multilingual attacks. DuoGuard uses a two-player RL-driven mechanism to generate harmful data in multiple languages and uses that to finetune a Llama3.2-1B model. PolyGuard collects an extensive dataset of 1.91M samples of harmful and benign datapoints in 17 languages and uses that to finetune a Qwen-2.5-7B-Instruct model.
@@
Relatively few works have tackled detecting harmful prompts in multimodal settings (see Table 1 and Table 2). Chi et al. (2024) propose LlamaGuard3-11B-Vision (a finetuned version of Llama-3-11B-Vision) for detecting unsafe inputs in images and texts. Du et al. (2024) and Helff et al. (2025) propose other approaches for the same task. OmniGuard achieves higher accuracy in detecting harmful images and prompts compared to these approaches, and to the best of our knowledge is the first guard model for harmful audio inputs.

### v2

@@ 3.2.3 Setup for audio attacks
Following Section˜3.1, we build a classifier that takes as input embeddings from layer 21 of Llama-Omni-8B. We train the classifier on the English portion of the training data we use for the text setting, by using a text-to-speech model to convert the text into audio. We use the open-source Kokoro model as the text-to-speech model.
@@
We are unaware of any existing models for detecting harmful audio input. The most relevant approach, SpeechGuard (peri-etal-2024-speechguard) adds noise as a defense against potentially harmful audio inputs but does not directly classify harmfulness. To contextualize our results for audio benchmarks, we compare performance to guard models that directly classify the raw text present in the audio (OmniGuard and LlamaGuard3).
@@
We use the two audio benchmarks (see Table˜1 bottom row). We also evaluate on several text jailbreak benchmarks using Kokoro to convert them from text to speech: HB, FQ, Simple Safety Tests, SaladB, and TJS. We use Kokoro for generating text-to-speech versions.
@@ 4 Results
Table 3: Accuracy of detecting harmful prompts for text attack benchmarks that are (A) multilingual benchmarks, (B) English translated to 73 languages, and (C) English translated to languages not seen at training time. In all settings, OmniGuard achieves the highest performance. Table˜B1 further stratifies these results by high-resource, low-resource, and cipher languages.
@@
Table˜3 compares the accuracy of detecting harmful prompts for text benchmarks. Table˜3(A) shows results for multilingual benchmarks, where OmniGuard achieves the highest accuracy (86.36%) compared to the baselines, and achieves new state-of-the-art performance for 3 benchmarks: MultiJail, RTP-LX, and AyaRedTeaming. The strongest baseline is Polyguard, which yields an average accuracy of 83.19%, despite being trained on a much larger dataset (1.91M examples for Polyguard versus 103K examples for OmniGuard). In benchmarks that were translated from English to various other languages, including cipher languages, we again see that OmniGuard achieves the highest accuracy (Table˜3(B)). Finally, Table˜3(C) shows that OmniGuard outperforms finetuned versions of DuoGuard and Polyguard on unseen languages, demonstrating that OmniGuard can outperform methods that were trained specifically for multilingual harmfulness classification.
@@
Table˜4 shows the accuracy of detecting harmful image and text prompts for (A) pairs consisting of images and text queries, where either the image or both the image and query can be harmful and (B) typographic images with various encryptions. OmniGuard achieves the highest performance for both sets of benchmarks (95.44% and 79.76%) while being trained using only about 3500 image-query pairs (compared to about 5500 datapoints used by LlavaGuard). The only benchmark where OmniGuard fails to detect harmful prompts is MML Base64, which consists of typographed images of prompts encrypted using base64 encoding.
@@
Table˜5 shows the accuracy of detecting harmful audio prompts. OmniGuard detects harmful audio input with high accuracy across all benchmarks. As we are not aware of any existing defenses for audio jailbreaks, we compare against OmniGuard and LlamaGuard3’s accuracy in detecting harmful prompts when the same inputs are provided in English text. The accuracy OmniGuard achieves in detecting harmful audio inputs is similar to or higher than its performance for detecting harmful text inputs.
@@
We also evaluate the accuracy of OmniGuard and baselines in adapting to out-of-distribution code attacks given very few samples. In this setting, some prior work has speculated that guard models may be very data efficient, as they can make use of few-shot examples in-context (inan2023-llama-guardrails). However, we find that baseline guard models generally struggle to rapidly adapt to this setting given few-shot examples (Figure˜3). [^2] Note that we omit baseline guard models that achieve 90% accuracy or greater without any few-shot examples, as their training data likely explicitly includes code attacks. In contrast, OmniGuard is able to rapidly achieve close to 100% accuracy for all three benchmarks by updating its lightweight parameters using less than five examples.
@@ 5 Analysis
We perform ablation experiments to determine the effect of selecting the appropriate layer for training the OmniGuard classifier. For the text-only model, we compare the U-Score-selected layer (57) to 3 other layers (layer 10, layer 75, and the last layer) when used for a set of toxicity prediction tasks. Table˜6 shows that the representations from the layer with the highest U-Score result in significantly better harmfulness classification accuracy, improving between 5% and 14% compared to the other layers. We show ablation over more layers in Table˜B2.
@@
Table 6: OmniGuard’s accuracy of detecting harmful prompts when trained using representations from different model layers.
@@
[…]
@@
[…]
Performance comparison across base LLMs
We compare OmniGuard’s accuracy when using different base LLMs in Table˜B3. We trained the classifiers on the layers with the best U-scores for each model. We find that the average accuracy for the moderator model trained using smaller LLMs is lower than the moderator model trained using the larger Llama3.3-70B-Instruct model.
@@
We now analyze the harmfulness classification accuracy of OmniGuard by language, and compare it to the underlying LLM’s sentiment classification accuracy for the same language (Fig.˜4). We measure harmfulness classification accuracy using OmniGuard on all the datasets in Table˜3 and sentiment classification accuracy using Llama3.3-70B-Instruct with zero-shot prompting on 72 translated versions of the SST-2 dataset (translated to all the languages we consider).
We observe that the accuracies are generally correlated, indicating that OmniGuard is able to defend well in languages for which the LLM is more coherent/susceptible to attack. Unsurprisingly, the accuracies for natural languages are higher than the accuracies for cipher languages. Nevertheless, harmfulness classification accuracy can be fairly high, even when sentiment classification accuracy is near chance (50%).
@@ 6 Related Work
Several techniques have recently emerged to attack or jailbreak LLMs. Early techniques relied on manual effort and were very time-intensive shen2024-do-now-evaluating; andriushchenko2024-adaptive-attacks-best-asr-paper. Later techniques automated this process, e.g., zou2023-gcg-attacks; jones-another-gcg-attack; zhu2023-autodan-interpretable-gradient proposed gradient-based approaches to identify inputs to jailbreak LLMs with white-box access. Another set of techniques start from a set of human written prompts and modify them using approaches like genetic algorithms (liu2024-autodan; lapid2024-opensesame-another-genetic; li2024-semanticmirror-another-genetic), fuzzing (yu2024-gpt-fuzzer-jailbreak), or reinforcement learning (chen2024-rl-jailbreak) to automatically produce prompts for jailbreaking. Another set of techniques, use a helper LLM to generate prompts that attack a target LLM (chao2024-PAIR-method; ding-2024-nested-attacks; mehrotra2024-tree-attacks-jailbreak). Finally, wei2024-icl-jailbreak1; wang2023-icl-jailbreak2; anthropic-many-shot-jailbreak; pernisi2024-many-shot-jailbreak-italian use simple in-context demonstrations to jailbreak the models by overcoming its safety training and russinovich2024-multi-turn-crescendo-attack; scale-ai-multi-turn-dataset propose using multi-turn dialogues to jailbreak models.
@@
Most of the aforementioned jailbreak techniques focus on attacks in English, against which significant defense exists both at the model and system level. To tackle this, a novel set of techniques have emerged that attack models using inputs in various languages or obfuscations that are able to bypass the safety guardrails. deng2024-multilingual-attack-firstpaper; yong2024-low-resource-languages-jailbreak-gpt4; wang2024-multilingualsafety; dan-roth-paper-guard-llms-on-multilingual-data; yoo2024-switch-between-languages; upadhayay2024-sandwich-attack-multilanguagemixture; song2024-multilingual-attacks demonstrated that attacking models using mid and low resources languages led to higher attack success rates, compared to the case of attacking the model in high-resource languages like English.
Going beyond natural languages, a newer set of works propose using cipher characters or languages to evade the safety filters, e.g., jin2024-JAMBench propose interspersing cipher characters in between text, jiang2024-ascii-jailbreak propose replacing the unsafe words with their ASCII art versions, and yuan2024-gpt-cipher-attack propose prompting models in cipher languages like Morse, Atbash, Caesar.
@@
Using modalities apart from text aims to explore a completely new attack surface, like images or audios. Several recent works have shown that MLLMs remain vulnerable to being jailbroken when prompted with images or audios that have a harmful query (the same harmful query in text would be easily detected as harmful). mm-safetybench-image-jailbreak show that using a prompt with a correlated image, e.g., using an image of a bomb when asking the model to answer the question: How to make a bomb? is more likely to jailbreak a model than when using an uncorrelated image. vlsbench-dataset-image-jailbreak argue that providing a harmful image with a benign query (see Figure˜1) further increases the potential of jailbreaking the model. figstep-dataset-image-jailbreak and mml-safebench-image-jailbreak demonstrate jailbreaking models by simply typographically embedding harmful queries in an image.
@@
Safety moderation in LLMs broadly fits into two categories: intrinsic and extrinsic. Intrinsic mechanisms include finetuning or RLHF training on an LLM (bianchi2024-finetuning-for-safety; chen2024-finetuning-for-safety; yuan2024-refuse-feel-unsafe; rlhf-paper1; rlhf-paper2; dai2023-safe-rlhf). Extrinsic safety mechanisms utilize external models to detect harmful inputs and responses; these models can either be simple filters or use guard models. jain2023-perplexity-defense1; alon2023-perplexity-defense2; hu2024-perplexity-defense3 propose using perplexity filtering for detecting harmful prompts. Guard models defend LLMs by training separate LLMs to detect harmful text (see Table˜2). Separately, a line of work has introduced interpretability methods to transparently expose safety concerns in LLMs (bereska2024mechanistic; singh2023augmenting; arditi2024refusal; benara2024crafting).
A few other works also use internal model representations to defend against harmful inputs. However, they defer from our work: OmniGuard is a standalone content safety classification model while the other approaches like Jailbreak Antidote (shen2025jailbreak-antidote) and AdaSteer (zhao2025-ada-steer) directly change the internal representations to defend against harmful inputs. Therefore, OmniGuard is similar to other safety classification models like LlamaGuard and WildGuard. Additionally, OmniGuard can detect harmful prompts in multiple languages (both natural and cipher) and multiple modalities using the same method while the other approaches only defend against attacks in English text. OmniGuard is also extremely efficient in producing safety predictions while the other approaches take as much time as the inference of the underlying LLM, which can typically take several seconds to minutes.
@@
DuoGuard (duo-guard-multilingual) and PolyGuard (polyguard-model) are the two previous guard models that were specifically trained to defend against multilingual attacks. DuoGuard uses a two-player RL-driven mechanism to generate harmful data in multiple languages and uses that to finetune a Llama3.2-1B model. PolyGuard collects an extensive dataset of 1.91M samples of harmful and benign datapoints in 17 languages and uses that to finetune a Qwen-2.5-7B-Instruct model.
@@
Relatively few works have tackled detecting harmful prompts in multimodal settings (see Table˜1 and Table˜2). chi2024-llamaguard3-vision propose LlamaGuard3-11B-Vision (a finetuned version of Llama-3-11B-Vision) for detecting unsafe inputs in images and texts. du2024-vlmguard and helff2025-llavaguard propose other approaches for the same task. OmniGuard achieves higher accuracy in detecting harmful images and prompts compared to these approaches, and to the best of our knowledge is the first guard model for harmful audio inputs.
@@ Appendix B Datasets and models
Appendix C Experimental Details of Filtering of Wikitext and its Translation
Thai Tox Ukr Tox TJS ToxText Avg.
Layer 10 62.1 65.5 66.95 61.89 64.42
Layer 55 67.4 73.0 76.91 74.96 73.06
Layer 56 66.8 71.5 74.54 71.82 71.17
Selected layer 57 68.7 73.1 76.8 75.07 73.40
Layer 58 66.6 73.2 74.92 72.63 71.84
Layer 59 67.3 73.3 76.44 74.22 72.82
Layer 60 67.5 72.6 76.43 74.46 72.75
Layer 61 67.8 70.9 74.77 72.76 71.56
Layer 62 66.1 72.3 74.78 72.83 71.50
Layer 75 65.2 66.4 70.72 65.79 68.26
Last Layer 63.1 51.2 61.33 56.76 59.05
Table B2: OmniGuard’s accuracy of detecting harmful prompts when trained using representations from different model layers.
Appendix D OmniGuard’s Performance With Different Base LLMs
Models HarmBench FQ SimpleST SaladB TJS ToxText AdvBench Avg.
Llama3.1-8B-Instruct 81.16 88.72 88.78 87.35 71.32 69.20 88.62 82.16
Gemma-3-4B-Instruct 78.03 88.23 93.86 82.09 53.50 51.45 89.07 76.60
Qwen-3-4B-Instruct 82.06 86.25 80.99 82.15 58.83 57.70 84.80 76.11
Olmo-2-7B-Instruct 80.57 88.86 92.14 88.68 70.65 64.79 88.24 81.99
Mistral-8B-Instruct 84.27 87.23 89.82 88.08 72.49 68.71 89.07 82.81
Llama3.3-70B-Instruct 89.13 89.57 89.62 87.30 76.68 75.07 86.59 84.85
Table B3: OmniGuard’s accuracy of detecting harmful prompts when paired with different underlying LLMs across multiple benchmarks.